# Broken intrinsic symmetry induced magnon-magnon coupling in synthetic ferrimagnets

### v1

@@ IV Summary
In summary, we demonstrated strong coupling between acoustic and optical modes in an in-plane geometry through broken intrinsic symmetry in a synthetic ferrimagnetic structure comprising of CoFe(10 nm)/Ru($x$ nm)/NiFe(10 nm). Micromagnetic simulations using Mumax3 reveal that both the gap field and the size of the gap depend on the quadratic interaction $J_{\mathrm{q}}$, while they are almost independent of the biquadratic interaction, $J_{\mathrm{bq}}$. Furthermore, the simulations allow us to unambiguously distinguish between the optical and acoustic modes observed in the experiment, and the phase analysis shows that the modes are indeed hybridized at the degeneracy field. Our study introduces intrinsic symmetry-breaking as a mechanism to tune magnon-magnon interaction in an in-plane geometry, providing intriguing opportunities for engineered magnonic hybrid systems.
Acknowledgements
This research was supported by NSF through the University of Delaware Materials Research Science and Engineering Center, DMR-2011824. The authors acknowledge the use of facilities and instrumentation supported by NSF through the University of Delaware Materials Research Science and Engineering Center, DMR-2011824.”
The data supporting this study’s findings are available upon reasonable request.

### v2

@@ IV Summary
In summary, we demonstrated strong coupling between acoustic and optical modes in a synthetic ferrimagnet, CoFe(10 nm)/Ru($x$ nm)/NiFe(10 nm), stabilized by broken intrinsic symmetry in an in-plane geometry. Macrospin model and micromagnetic simulations reveal that both the gap field and gap size depend sensitively on the quadratic and biquadratic exchange interactions, $J_{\mathrm{q}}$ and $J_{\mathrm{bq}}$. While $J_{\mathrm{q}}$ and $J_{\mathrm{bq}}$ can in principle be extracted from magnetostatics via Eqs. (3a) and (3b), this approach lacks precision and fails for $J_{\mathrm{q}}<2J_{\mathrm{bq}}$ where $H_{0,\mathrm{afm}}$ does not exist. By contrast, the macrospin model developed here provides a robust route to parameter extraction by fitting directly to spin-dynamics spectra. Moreover, micromagnetic simulations enable clear distinction between acoustic and optical modes and, through phase analysis, reveal hybridization at the degeneracy field. Our micromagnetic framework, incorporating both quadratic and biquadratic interactions, is broadly applicable to unconventional geometries, including artificial spin-ice structures Sultana et al. (2025); Dion et al. (2024), where the macrospin models break down. Furthermore, our study introduces intrinsic symmetry-breaking as a mechanism to tune magnon-magnon coupling in an in-plane geometry, which offers promising prospects for the development of reconfigurable magnonic devices, such as tunable filters, non-reciprocal components, and coherent magnon-based logic elements. The ability to engineer and control hybridized modes through intrinsic symmetry-breaking can significantly enhance the design flexibility of future on-chip spintronic and magnonic circuits.
Acknowledgments
This research was supported by NSF through the University of Delaware Materials Research Science and Engineering Center, DMR-2011824. The authors acknowledge the use of facilities and instrumentation supported by NSF through the University of Delaware Materials Research Science and Engineering Center, DMR-2011824.
The data supporting this study’s findings are at [?].